# DFREC: DeepFake Identity Recovery Based on Identity-aware Masked Autoencoder

### v1

@@ Abstract
Article submission, IEEE, IEEEtran, journal, , paper, template, typesetting.
@@ I Introduction
The emergence of deepfake technology, particularly its face-swapping capabilities, has ushered in an unprecedented era of digital media manipulation. It enables the seamless substitution of one individual’s face with another’s. Since its inception, deepfake has rapidly evolved into tools for disseminating misinformation, identity theft, and fraud. The growing ethical and legal concerns on the widespread dissemination of deepfake content have underscored the urgency of developing robust deepfake forensic tools. This is crucial not only for safeguarding individuals’ rights but also upholding the information integrity in the digital age.
Most of the existing deepfake forensic methods focus on improving the generalization detection performance across various kinds of forgery images [1, 2, 3, 4, 5, 6]. Despite their commendable performance, detection-based algorithms are limited to providing classification or localization results. They lack the capability to provide a corresponding chain of evidence to match the detection result. In other words, they cannot trace which are the two faces that have been swapped. The inability to trace the origin of deepfake content can delay timely action and reduce the effectiveness of incident response plan to mitigate potential damage. More importantly, as the onus of proof is borne by the victims with source identities, an accurate recovery of the source and target identities could provide an irrefutable evidence of manipulation. The recovered faces can help to protect the victim from potential hazards and in collecting and identifying evidences for pursuing legal action against the defendant responsible for creating and disseminating deepfake content with malicious intent.
@@
Figure 1: (a) The training process of deepfake face swapping models. The identity loss is used to increase the similarity between the forgery and source face identities, and the attribute loss is used to ensure the similarity between the forgery and target face attributes. (b) The probability density functions of the identity similarity (IDSim) between the forgery and source, forgery and target, and forgery and other unrelated faces.
@@
Motivated by this forensic investigation gap, we propose to approach the deepfake forensic task from the perspective of pristine face recovery. While deepfake algorithms have evolved through various iterations, their core operation involves transferring the source identity onto the target face while retaining the attributes of the target face (such as expression, pose, lighting, etc.). This synthesis process inadvertently blends the identities of both faces, resulting in a composite image that incorporates features from both identities [7]. As illustrated in Figure 1, the identity similarity between the forgery face and the target face is closely resembling the similarity between the forgery face and the source face. This observation suggests that the forgery face encompasses a portion of the target face information. Building on this insight, we propose DFREC, a DeepFake Identity Recovery scheme based on Identity-aware Masked autoencoder, which is designed to recover both the source and target faces. The main contributions of this paper are outlined as follows:
@@
  • We propose a digital forensic aid DFREC to restore the identities of both target and source faces. DFREC accomplishes this task through ($i$) designing an Identity Segmentation Module (ISM) to segment information related to the source and target faces respectively, ($ii$) constructing a Source Identity Recovery Module (SIRM) to recover the source face using the segmented source information, and ($iii$) designing a Target Identity Recovery Module (TIRM) to recover the target face using the background information and identity features extracted by the SIRM. Through identity segmentation and recovery, DFREC is able to restore the pristine source and target faces with high fidelity for intuitive interpretation.
@@ II-A Deepfake Face Swapping
Face swapping is a representative algorithm in deepfake technology. It can be categorized into source-oriented and target-oriented methods. Source-oriented methods primarily involve warping the source face to align with the target’s pose and expression before it is seamlessly merged into the target frame [8, 9, 10]. Target-oriented methods, on the other hand, leverage generative models to encode target features and synthesize the swapped face [11, 12, 13, 14, 15, 16]. It is worth noting that these face swapping algorithms produce high-fidelity forgery face by preserving target’s face attributes (such as expression and pose). Due to the inherent difficulty in decoupling identity features from attribute features, the extracted attribute features often exhibit a certain degree of correlation with identity [7]. This inherent correlation often leads to a fusion of source and target identities. Inspired by this insight, our proposed DFREC scheme extracts latent identity features from forgery images to recover the source and target faces, thereby enabling the tracing of both source and target identities.
@@ II-B Deepfake Forgery Detection
Deepfake forgery detection predominantly utilizes CNN-based methodologies, and focus more on the generalization performance in recent years. Generalization through diverse strategies, including data augmentation [17, 1, 18], feature consistency analysis [19, 20], and frequency domain anomalies  [21, 22] have been explored in recent years to detect fine-grained photorealistic face forgery. Despite their high detection accuracy, these forensic tools cannot trace the origin of deepfake content, which prevents the victims from pursuing legal action and initiating timely incident response to minimize damage.
@@ II-C Deepfake Recovery
Different from detection-based methods, deepfake recovery methods aim to reconstruct the original target face. To improve the identity-relevant deepfake detection and interpretability, RECCE [23] proposed a reconstruction learning approach to discern compact real data representations. Delocate [24] utilized a Masked Autoencoder to reconstruct original facial components to assist in deepfake detection. Additionally, DFI [25, 26] proposed to trace the original target face for interpretable deepfake detection. However, RECCE and Delocate are limited to generating distorted facial images and cannot effectively recover the target identity from the forgery face. DFI cannot avoid the disturbance of source identities on the recovered target faces. To tackle these limitations, we first segment the two types of information within forgery faces to mitigate their interference, and then integrate the background information with the implicit target identity features to recover the deepfake identity.
@@ III The Proposed Method
We name our proposed deepfake identity recovery scheme DFREC. As illustrated in Figure 2, DFREC consists of three components: an Identity Segmentation Module (ISM), a Source Identity Reconstruction Module (SIRM), and a Target Identity Reconstruction Module (TIRM). The ISM classifies each pixel of an input image according to its probability of affiliation with the source face. By taking the dot product between the input image and the output probability map, the segmented source face information is produced. The SIRM reconstructs the source face and extracts the target identity features from the decomposed source face information simultaneously. The TIRM masks the image using the segmentation map and recovers the target face by combining the background information with the target identity features. A quadruple loss that combines Identity Loss, Perceptual Loss, Attribute Loss and Patch Recovery Loss is used to guide the face recovery to achieve state-of-the-art (SOTA) recovery quality of the source and target faces.
III-A Identity Segmentation Module (ISM)
@@
The ISM is designed to segment source and target face information embedded in a face-swapping image. Specifically, we employ the DeepLabv3 architecture based on atrous convolution, which consists of a feature extractor, a segmentation head and a classification head. The classification head determines if an input facial image is a pristine or a forgery face while the segmentation head produces a probabilistic map for the extraction of source face information. We use pretrained EfficientNet-B0 as in [27] to extract both low-level and high-level features. The high-level features are then directed into the classification head for forgery classification. Meantime, the high-level features are supplied to the segmentation head for up-sampling and combined with the low-level features to generate the segmentation map. As illustrated in Figure 2, it takes the image $I$ as input and generates a segmentation map $m_{s}$. This map is used to produce a segmented image $I_{src}^{seg}$ for source face recovery as well as guide the MAE to iteratively mask source-related patches for the target face recovery. It is worth noting that we refrain from utilizing any loss function to constrain the segmentation process, as it will autonomously compute the necessary information driven by subsequent tasks.
III-B Source Identity Recovery Module (SIRM)
@@
The SIRM is designed to recover the source face and extract the latent target identity features using the segmented $I_{src}^{seg}$. Specifically, we construct a UNet-like model for source face recovery. As depicted in Figure 2, the SIRM is made up of an encoder and a decoder. The encoder is responsible for identity feature extraction. The extracted identity features are then utilized by the decoder for the source face recovery. Due to the fusion of two identities, features extracted by the encoder may inevitably contain some target face features. This may detrimentally mislead the source face recovery. To address this problem, the features extracted by the encoder are divided into two parts: source identity feature $f_{s}$ and target identity feature $f_{t}$. The separated $f_{s}$ is independently exploited by the decoder for the recovery of the source face, and the separated $f_{t}$ is utilized for subsequent target face recovery. An attention mechanism is then introduced into the skip connection to preserve features pertinent to the source face. Finally, the source face is reconstructed from the segmented source information.
III-C Target Identity Recovery Module (TIRM)
@@
The TIRM is devised to recover the target face image. Given that deepfake face swapping algorithms aim to eliminate as much of the target identity from the swapped image as possible, restoring the target face poses more challenges than restoring the source face. To tackle this issue, we introduce an Identity-aware Masked Autoencoder (MAE) to combine the background information and the predicted target identity features to recover the target face. As illustrated in Figure 2, the TIRM comprises two stages: Semantic-Guided Masking and Identity Recovery.
@@
Semantic-Guided Masking: This stage masks a fixed proportion of patches under the guidance of the predicted segmentation map $m_{s}$. Traditional MAE-based methods typically mask image patches randomly, and then use visible patches to recover the masked patches. However, most of the forged regions are concentrated around the facial area. Random masking has high propensity to preserve some strong forged facial components that can mislead the target face recovery process. We mitigate this problem by using the predicted $m_{s}$ to select image patches to be masked. As demonstrated in Algorithm 1, we first segment $m_{s}$ into $n$ patches, each of size $\rho$. Each patch is weighted by its average value. For patches with weights less than 0.5, their weights are replaced by random numbers between 0 and 0.1. Then, the patch weights are sorted in descending order. The top $m$ highest weight patches are selected, where the number of patches to be masked $m$ is determined by the mask ratio $\lambda$. This strategy gradually shifts the patch masking from random to source-related areas as the accuracy of the predicted segmentation map improves. Eventually, all patches related to the source face will be masked and inpainted by using other visible patches to restore the target face.
@@
0:  Indices of the masked and kept patches.
@@
4:  Generate random vector: $\nu\leftarrow\text{RandVector}(n)$;
@@
Identity Recovery: The aim of this step is to restore the masked patches to the corresponding patches in the target face. We utilize the background contextual information and the latent target identity information to achieve this objective. Background context information refers to the unmasked patches within the input image, where structural and color correlations among various image components exist. These correlations between pixels or regions can assist in predicting partial information about the masked areas from the unmasked patches. The target identity information is the latent features $f_{t}$ extracted by the SIRM to guide the reconstruction of the target face. To synergize the background context information with the latent target identity information, we construct an Identity Fusion Decoder in the MAE. The target identity features and visible patch embeddings are concatenated to predict the masked patches. Its principal operations are delineated as follows.
@@ IV Experiments
MAT 21.01 0.4372 19.23 0.3874 26.98 0.4026 42.06 0.3928 34.64 0.4203 35.98 0.3506 9.82 0.2797
Repaint 21.44 0.3832 19.49 0.3366 27.52 0.3639 41.18 0.3874 33.97 0.3893 35.11 0.3423 38.97 0.3946
RECCE 87.28 0.3755 65.72 0.4400 89.33 0.4086 90.36 0.2547 115.35 0.2957 110.69 0.2313 81.82 0.2668
Delocate 18.16 0.4749 37.22 0.4551 135.93 0.3858 40.75 0.4201 39.90 0.4650 41.76 0.3466 36.88 0.3064
DFI 42.17 0.4706 30.39 0.5207 37.03 0.4860 22.19 0.4137 22.20 0.4346 29.30 0.3442 29.86 0.3805
@@
IV-A Experimental Settings
IV-A1 Datasets
We utilize three deepfake datasets to train and validate the performance of our scheme. For Faceforensics++ [34], we select two existing forgeries, namely DeepFake (DF) and FaceShifter (FShi). Additionally, we generate another forgery set using the LCR algorithm [35]. The three types of forgery images are then splitted into their corresponding training, validation, and testing images according to the splitting strategy in [34]. CelebaMegaFS [35] incorporates three deepfake methods: IDInjection (IDI), FTM, and LCR, resulting in 30,038 FTM images, 30,441 IDInjection images, and 30,010 LCR images. FFHQ-E4S [16] is constructed based on the FFHQ dataset [36]. We randomly select 10,000 images from FFHQ and use E4S [16] to generate 10 swapped images per original image. For CelebaMegaFS and FFHQ-E4S, each dataset is divided into the training, validation, and test sets according to an 8:1:1 ratio. In the Faceforensics++ and FFHQ-E4S datasets, the identities in the training and testing sets are non-overlapping. In CelebaMegaFS, about 70% of the test identities overlap with the training set.
IV-A2 Compared approaches
We compare DFREC with five SOTA methods. MAT [37] is a mask-aware transformer for large hole image inpainting, which combines transformers and convolutions to reconstruct high resolution images. RePaint [38] is a face inpainting method which uses a pretrained unconditional Denoising Diffusion Probabilistic Model (DDPM) as the generative prior. RECCE [23] is a reconstruction-based deepfake forensic method, which exploits the reconstruction difference as guidance of forgery traces. Delocate [24] utilizes an MAE to learn the distribution of pristine faces for the recovery of manipulated faces. DFI [26] is a deepfake inversion method that learns the reverse mapping from the forgery face to the original face.

### v2

@@ Abstract
Deepfake, identity recovery, mask autoencoder, image forensic.
@@ I Introduction
The emergence of deepfake technology, particularly its face-swapping capabilities, has ushered in an unprecedented era of digital media manipulation. Since its inception, deepfake has rapidly evolved into tools for disseminating misinformation, identity theft, and fraud. The growing ethical and legal concerns on the widespread dissemination of deepfake content have underscored the urgency of developing robust deepfake forensic tools. This is crucial not only for safeguarding individuals’ rights but also upholding the information integrity in the digital age.
(a) Deepfake Face Swapping Model Training
(b) Identity similarity between forgery and target/source faces
@@
Figure 1: (a) The training process of deepfake face swapping models. The identity loss is used to increase the similarity between the forgery and source face identities, and the attribute loss is used to ensure the similarity between the forgery and target face attributes. (b) The probability density functions of the identity similarity (IDSim) between the forgery and source, forgery and target, and forgery and other unrelated faces. The results are calculated on the five types of forgeries on the FaceForensics++ and CelebaMegaFS datasets.
Most of the existing deepfake forensic methods focus on improving the generalization detection performance across various kinds of forgery images [1, 2, 3, 4, 5, 6, 7]. While these methods demonstrate impressive performance in classification and localization tasks, they are inherently limited in their ability to provide comprehensive forensic evidence. Specifically, they fail to establish a coherent chain of evidence linking the detection results to the identities involved in the manipulation. In other words, they cannot reliably trace the source identity (i.e., the person whose facial features are transferred) and the target identity (i.e., the individual whose face appears in the manipulated content). The inability to trace the origin of deepfake content can delay verifying the origin of deepfake content and mitigating potential damage. More importantly, as the onus of proof is borne by the victims with source identities, an accurate recovery of the source and target identities could help to collect and identify evidences for pursuing legal action against the defendant responsible for creating and disseminating deepfake content with malicious intent.
@@
Motivated by this forensic investigation gap, we propose to approach the deepfake forensic task from the perspective of pristine face recovery. While deepfake algorithms have evolved through various iterations, their core operation involves transferring the source identity onto the target face while retaining the attributes of the target face (such as expression, pose, lighting, etc.). This synthesis process blends the identities of both faces, resulting in a composite image that incorporates features from both identities [8]. As illustrated in Fig. 1, the identity similarity between the forgery face and the target face is closely resembling the similarity between the forgery face and the source face. This observation suggests that the forgery face encompasses a portion of the target face information. Building on this insight, we propose DFREC, a DeepFake Identity Recovery scheme based on Identity-aware Masked autoencoder, which is designed to recover both the source and target faces. The main contributions of this paper are outlined as follows:
@@
  • We propose a digital forensic aid DFREC to restore the identities of both target and source faces. DFREC accomplishes this task through ($i$) designing an Identity Segmentation Module (ISM) to segment information related to the source and target faces respectively, ($ii$) constructing a Source Identity Recovery Module (SIRM) to recover the source face using the segmented source information, and ($iii$) designing a Target Identity Recovery Module (TIRM) to recover the target face using the background information and latent target identity features extracted by the SIRM. Through identity segmentation and recovery, DFREC is able to restore the pristine source and target faces with high fidelity for intuitive interpretation.
@@ II Related Works
In this section, we first review existing deepfake face-swapping algorithms, then discuss current detection methods, and finally introduce the development of deepfake recovery techniques.
@@ II-A Deepfake Face Swapping
Face swapping is a representative algorithm in deepfake technology, which leverages generative models to encode target features and synthesize the swapped face. Convolutional neural networks (CNNs) and generative adversarial networks (GANs) are often employed to achieve identity-attribute encoding [9, 10, 11, 12]. Gu et al. [9] proposed to leverage conditional GANs to generate specific swapping faces. This framework learns feature embeddings for each facial component independently, enabling precise and customizable local face editing. To further improve the visual quality of the swapped faces, Li et al. proposed FaceShifter [10], which employs a U-Net architecture to capture the target face’s attributes. It incorporates an Adaptive Embedding Integration Network to blend source facial features with target attributes, producing photorealistic face-swapping results. Similarly, SimSwap [11] introduces a Weak Feature Matching Loss to optimize generated faces. It exploits both shallow and deep discriminator features for better consistency and fidelity. Based on these advancements, Zhu et al. [12] proposed MegaFS, a novel approach that operates in an extended latent space to generate megapixel-level swapped images. It achieves superior identity swapping while preserving high-resolution facial attributes. Recent advances [13, 14, 15] further enhance realism and preserve the geometry and texture priors, as well as other non-facial attributes in swapped images.
It is worth noting that these advanced face swapping algorithms produce high-fidelity forgery face identity by preserving the target’s face attributes (such as expression and pose). Due to the inherent difficulty in decoupling identity features from attribute features, the extracted attribute features often exhibit a certain degree of correlation with identity [8]. This inherent correlation often leads to a fusion of source and target identities. Inspired by this insight, our proposed DFREC scheme extracts latent identity features from forgery images to recover the source and target faces, thereby enabling the tracing of both source and target identities.
@@ II-B Deepfake Forgery Detection
Deepfake forgery detection predominantly utilizes CNN-based methodologies, and focus more on the generalization performance in recent years [16]. Generalization through diverse strategies, including data augmentation [17, 1, 18] and feature engineering [19, 20, 21, 22, 23], have been explored to detect fine-grained photorealistic face forgery. In the realm of data augmentation, Li et al. [17] proposed a method that selects two facial images based on the similarity of their landmarks, generating blended faces by applying randomly created masks on facial regions. This strategy significantly improved the generalization performance of detection models. Shiohara et al. [1] further introduced simulation based inference (SBI) to simulate challenging forgery images by transforming local regions of pristine images. Building on this foundation, Nguyen et al. [18] proposed LAANet, which calculates vulnerable points in the SBI framework and employs an attention mechanism to enhance sensitivity to these regions, achieving state-of-the-art generalization accuracy. In terms of feature engineering, Zhao et al. [19] enhanced forgery detection by constraining the correlation between features of authentic and forged regions to improve the model’s perception of manipulated content. Ni et al. [20] developed a method that enforces feature consistency across diverse augmentation operations, which facilitates the learning of stable and reliable representations. More recently, Yan et al. [21] proposed to expand the diversity of feature representations by transforming latent space, which effectively improves the detection performance against previously unseen forgeries. Despite their high detection accuracy, these forensic tools cannot trace the origin of deepfake content, which hinders the victims from pursuing legal action and limits their applicability in judicial contexts.
@@ II-C Deepfake Recovery
Different from detection-based methods, deepfake recovery methods aim to reconstruct the original target face. Existing studies primarily leverage face reconstruction to facilitate deepfake detection. To improve the identity-relevant deepfake detection and interpretability, RECCE [24] proposed a reconstruction learning approach to discern compact real data representations. They trained an autoencoder with real face images to reconstruct faces and then utilized the face reconstruction error to guide the training of deepfake models. Delocate [25], on the other hand, randomly occluded local regions of faces and employed a Masked Autoencoder to reconstruct the original facial components, thereby assisting deepfake detection. Additionally, DFI [26, 27] proposed to trace the original target face by decoupling identity and attributes. However, RECCE and Delocate are limited to generating distorted facial images and cannot effectively recover the target identity from the forgery face. DFI cannot avoid the disturbance of source identities on the recovered target faces. To tackle these limitations, we first segment the two types of information within forgery faces to mitigate their interference, and then integrate the background information with the implicit target identity features to recover the deepfake identity.
@@ III-A Overview
We name our proposed deepfake identity recovery scheme DFREC. As illustrated in Fig. 2, DFREC consists of three components: an Identity Segmentation Module (ISM), a Source Identity Reconstruction Module (SIRM), and a Target Identity Reconstruction Module (TIRM). The ISM classifies each pixel of an input image according to its probability of affiliation with the source face. By taking the dot product between the input image and the output probability map, the segmented source face information is produced. The SIRM reconstructs the source face and extracts the target identity features from the decomposed source face information simultaneously. The TIRM masks the image using the segmentation map and recovers the target face by combining the background information with the target identity features using the Masked Autoencoder (MAE). A quadruple loss that combines Identity Loss, Perceptual Loss, Attribute Loss and Patch Recovery Loss is used to guide the face recovery to achieve state-of-the-art (SOTA) recovery quality of the source and target faces.
III-B Identity Segmentation Module (ISM)
@@
The ISM is designed to segment source and target face information embedded in a face-swapping image. Specifically, we employ the DeepLabv3 architecture based on atrous convolution, which consists of a feature extractor, a segmentation head and a classification head. The classification head determines if an input facial image is a pristine or a forgery face while the segmentation head produces a probabilistic map for the extraction of source face information. We use pretrained EfficientNet-B0 as in [28] to extract both low-level and high-level features. The high-level features are then directed into the classification head for forgery classification. Meantime, the high-level features are supplied to the segmentation head for up-sampling and combined with the low-level features to generate the segmentation map. As illustrated in Fig. 2, it takes the image $I$ as input and generates a segmentation map $m_{s}$. This map is used to produce a segmented image $I_{src}^{seg}$ for source face recovery as well as guide the MAE to iteratively mask source-related patches for the target face recovery. It is worth noting that we refrain from utilizing any loss function to constrain the segmentation process, as it will autonomously compute the necessary information driven by subsequent tasks.
III-C Source Identity Recovery Module (SIRM)
@@
The SIRM is designed to recover the source face and extract the latent target identity features using the segmented image $I_{src}^{seg}$. Specifically, we construct a UNet-like model for source face recovery. As depicted in Fig. 2, the SIRM is made up of an encoder and a decoder. The encoder is responsible for identity feature extraction. The extracted identity features are then utilized by the decoder for the source face recovery. Due to the fusion of two identities, features extracted by the encoder may inevitably contain some target face features. This may detrimentally mislead the source face recovery. To address this problem, the features extracted by the encoder are divided into two parts: source identity feature $f_{s}$ and target identity feature $f_{t}$. The separated $f_{s}$ is independently exploited by the decoder for the recovery of the source face, and the separated $f_{t}$ is utilized for subsequent target face recovery. An attention mechanism (represented by $\triangle$ in Fig. 2) is then introduced into the skip connection to preserve features pertinent to the source face. Finally, the source face is reconstructed from the segmented source information.
III-D Target Identity Recovery Module (TIRM)
@@
The TIRM is devised to recover the target face image. Given that deepfake face swapping algorithms aim to eliminate as much of the target identity from the swapped image as possible, restoring the target face poses more challenges than restoring the source face. To tackle this issue, we introduce an Identity-aware Masked Autoencoder (MAE) to combine the background information and the predicted target identity features to recover the target face. As illustrated in Fig. 2, the TIRM comprises two stages: Semantic-Guided Masking and Identity Recovery.
III-D1 Semantic-Guided Masking
@@
This stage masks a fixed proportion of patches under the guidance of the predicted segmentation map $m_{s}$. Traditional MAE-based methods typically mask image patches randomly, and then use visible patches to recover the masked patches. However, most of the forged regions are concentrated around the facial area. Random masking has high propensity to preserve some strong forged facial components that can mislead the target face recovery process. We mitigate this problem by using the predicted $m_{s}$ to select image patches to be masked. As demonstrated in Algorithm 1, we first segment $m_{s}$ into $n$ patches, each of size $\rho$. Each patch is weighted by its average value. For patches with weights less than 0.5, their weights are replaced by random numbers between 0 and 0.1. Then, the patch weights are sorted in descending order. The top $m$ highest weight patches are selected, where the number of patches to be masked $m$ is determined by the mask ratio $\lambda$. This strategy gradually shifts the patch masking from random to source-related areas as the accuracy of the predicted segmentation map improves. Eventually, all patches related to the source face will be masked and inpainted by using other visible patches to restore the target face.
@@
0:  Indices of the masked and unmasked patches.
@@
4:  Generate random vector of elements $\in[0,0.1]$: $\nu\leftarrow\text{RandVector}(n)$;
@@
III-D2 Identity Recovery
The aim of this step is to restore the masked patches to the corresponding patches in the target face. We utilize the background contextual information and the latent target identity information to achieve this objective. Background context information refers to the unmasked patches within the input image, where structural and color correlations among various image components exist. These correlations between pixels or regions can assist in predicting partial information about the masked areas from the unmasked patches. The target identity information is the latent features $f_{t}$ extracted by the SIRM to guide the reconstruction of the target face. To synergize the background context information with the latent target identity information, we construct an Identity Fusion Decoder in the MAE. The target identity features and visible patch embeddings are concatenated to predict the masked patches. Its principal operations are delineated as follows.
@@ IV Experiments
IV-A Experimental Settings
IV-A1 Datasets
We utilize three deepfake datasets to validate the performance of our scheme:
  1. -
    Faceforensics++ [35]: This dataset comprises 1000 segments of real videos, primarily centered around news broadcasts. We select two algorithms, namely DeepFake (DF) and FaceShifter (FShi), which contain a total of 2000 fake videos. Additionally, the real videos are processed using the LCR algorithm [12] to generate another set of 1000 fake videos. Following the splitting strategy outlined in [35], we obtain 100,000 training images, 20,000 validation images, and 20,000 testing images for each forgery method.
  2. -
    CelebaMegaFS [12]: This dataset is constructed from the Celeba-HQ dataset [36] for face swapping. It incorporates three deepfake methods: IDInjection (IDI), FTM, and LCR, resulting in 30,038 FTM images, 30,441 IDInjection images, and 30,010 LCR images. Each dataset is divided into training, validation, and test sets according to an 8:1:1 ratio.
  3. -
    FFHQ-E4S [14]: This dataset is constructed based on the FFHQ dataset [37]. We randomly select 10,000 images from FFHQ and use E4S [14] to generate 10 swapped images per original image. These images are distributed across training, validation, and test sets in an 8:1:1 ratio based on face identities.
IV-A2 Compared approaches
We compare DFREC with the following SOTA methods:
  1. -
    MAT [38]: It is a mask-aware transformer for large hole inpainting, which combines transformers and convolutions to reconstruct high resolution images.
  2. -
    RePaint [39]: It is a face inpainting method which uses a pretrained unconditional Denoising Diffusion Probabilistic Model (DDPM) as the generative prior.
  3. -
    RECCE [24]: It is a reconstruction-based deepfake detection method, which exploits the reconstruction difference as guidance of forgery traces.
  4. -
    Delocate [25]: It is a deepfake face recovery approach that utilizes MAE to learn the distribution of pristine faces for the recovery of manipulated faces.
  5. -
    DFI [27]: It is a deepfake inversion method that learns the reverse mapping from the forgery face to the original face.
@@ IV-A2 Compared approaches
MAT [38] 21.01 0.4372 19.23 0.3874 26.98 0.4026 42.06 0.3928 34.64 0.4203 35.98 0.3506 9.82 0.2797
Repaint [39] 21.44 0.3832 19.49 0.3366 27.52 0.3639 41.18 0.3874 33.97 0.3893 35.11 0.3423 38.97 0.3946
RECCE [24] 87.28 0.3755 65.72 0.4400 89.33 0.4086 90.36 0.2547 115.35 0.2957 110.69 0.2313 81.82 0.2668
Delocate [25] 18.16 0.4749 37.22 0.4551 135.93 0.3858 40.75 0.4201 39.90 0.4650 41.76 0.3466 36.88 0.3064
DFI [27] 42.17 0.4706 30.39 0.5207 37.03 0.4860 22.19 0.4137 22.20 0.4346 29.30 0.3442 29.86 0.3805